OpenStudy (anonymous):

2x + y + z = –7 x – 3y + 4z = –14 x – 2y – 3z = –11

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

Are you attemtping to solve for values of x y and z?

OpenStudy (anonymous):

@sammyrayepatton

OpenStudy (anonymous):

to explain which method I would use and how to solve it

OpenStudy (anonymous):

elimination method..

OpenStudy (anonymous):

yes, I just dont know how to do it, and it would be amazing if someone could help me!:/

OpenStudy (anonymous):

In the bottom two equations you have 1x in one equation and 1x in the other equation as well

OpenStudy (anonymous):

solve 2 eqns to remove a variable ,solve the new eqn with the 3rd eqn

OpenStudy (anonymous):

could you guys please show me how to do that?:/

OpenStudy (anonymous):

x-3y+4z=-14 - x-2y-3z= -11

OpenStudy (anonymous):

Use that so you can eliminate x from those two equations and proceed to solve the equation from there

OpenStudy (anonymous):

wait so what do I do with it, I'm trying to pass a class, and my grade is totally dependening on it, and I'm going to lose my mind I cant get this part of the lesson down..

OpenStudy (anonymous):

Alright, so when you subtract the bottom equation from the middle equation what do you get? ( The equation I just posted)

OpenStudy (anonymous):

x-5y+z=-25?

OpenStudy (anonymous):

Close, since you are subtracting Positive x from positive x, the x's would cancel out. Understand?

OpenStudy (anonymous):

5y+z=-25?